OpenStudy (blackops2_143):

What is the product in lowest terms? -6/11 x 3/4? A. -9/22 b. -6/13 c. 9/22 d. 6/13

OpenStudy (theprofessor35):

Multiply straight across -6 * 3 = ?

OpenStudy (theprofessor35):

-18/ 11*4 = -18/44

OpenStudy (theprofessor35):

Now simplify

OpenStudy (theprofessor35):

A. -9/22

OpenStudy (theprofessor35):

We are done! :)

OpenStudy (blackops2_143):

thank you!!